# Convergence of Fast Policy Iteration in Markov Games and Robust MDPs

### v1

@@ 1 Introduction
Markov Games (MG) (Kallenberg 2022) and Robust MDPs (RMDPs) (Iyengar 2005; Wiesemann, Kuhn, and Rustem 2013; Ho, Petrik, and Wiesemann 2022) are two important models that generalize Markov Decision Processes (MDPs) (Puterman 2005). Markov games can model strategic adversaries that can act to minimize the agent’s returns and are a common model in multi-agent reinforcement learning (Shou et al. 2022; Littman 1994). Similarly, RMDPs can model an adversarial nature that can perturb transition probabilities and rewards to minimize the agent’s returns and are useful when making decisions with imperfect data-driven models (Lobo et al. 2023; Behzadian et al. 2021). In recent years, MGs and RMDPs have seen an increasing number of applications in machine and reinforcement learning, which has motivated the study of efficient algorithms for solving them (Pérolat et al. 2016; Ho, Petrik, and Wiesemann 2021, 2022; Behzadian, Petrik, and Ho 2021; Kaufman and Schaefer 2013; Winnicki and Srikant 2023).
@@
Although basic algorithms, like value and policy iteration, adapt readily from MDPs to MGs and RMDPs, developing more efficient algorithms has been challenging. The efforts to adapt efficient optimistic policy iteration (OPI) algorithms, such as modified or fitted policy iteration, have been difficult. Many natural OPI algorithms proposed for MGs and RMDPs cycle among suboptimal policies, alternatively improving the minimization or maximization sides of the saddle point equilibrium. The lack of convergence is often counter-intuitive and has led to several incorrect convergence proofs in the literature (Condon 1993; Pérolat et al. 2016; Filar and Tolwinski 1991). The overarching reason is that OPI in MG and RMDPs do not monotonically improve the policy and its value function as in MDPs.
@@
We make two main contributions in this paper. First, we show that a fast OPI method proposed in Filar and Tolwinski (1991) can terminate with an arbitrarily suboptimal policy. This result confirms a suspicion, expressed in Pérolat et al. (2016), that there is a gap in the proof of the algorithm’s correctness.
@@
Second, we propose and analyze Residual Conditioned Policy Iteration (RCPI). RCPI is a new, simple approximate policy iteration algorithm for solving MGs and RMDPs that is guaranteed to converge to optimal policies. It builds on earlier efficient OPI algorithms (Filar and Tolwinski 1991; Pérolat et al. 2016; Ho, Petrik, and Wiesemann 2021; Winnicki and Srikant 2023) and combines them with an adaptive correction step. Our theoretical analysis shows that RCPI matches the worst-case computational complexity of value iteration. Our numerical results show that on a wide range of problems, RCPI outperforms other convergent algorithms by several orders of magnitude, even in moderately sized problems.
@@
In this paper, we restrict our focus to model-based algorithms for MGs and RMDPs. It is important to note that this setting differs from online algorithms for solving games and multi-agent reinforcement learning problems, such as in Zhang et al. (2022). Although some of the issues that need to be overcome in online and model-based solvers are similar, we leave the study of the exact relationship between online and model-based algorithms for future work.
@@
The remainder of the paper is organized as follows. Section˜2 positions our work in the context of prior algorithmic developments for MGs and RMDPs. Then, Section˜3 describes the formal framework for MGs and RMDPs. Section˜4 describes our first contribution, which is to show that an existing OPI algorithm (Filar and Tolwinski 1991) may fail with an arbitrarily suboptimal policy. Section˜5 describes our second and main contribution, the RCPI algorithm, along with its convergence rate and computational complexity analysis. Finally, our numerical results in Section˜6 compare RCPI with the state-of-the-art algorithms for solving MGs and RMDPs.
@@ 2 Prior Work: Solving MGs and RMDPs
In this section, we summarize prior efforts on developing OPI algorithms for MGs and RMDPs. We note that MG and RMDP communities have been largely separate, though the similarities between them have been noted and exploited previously (Iyengar 2005; Grand-Clément and Petrik 2024; Grand-Clément, Petrik, and Vieille 2025).
Value iteration is a simple convergent algorithm for solving MDPs, RMDPs, and MGs, but can be very slow in many practical settings (Puterman 2005). Many faster convergent algorithms for MDPs exist, such as policy iteration or modified policy iteration. Since the early days of MG (Condon 1993) and RMDPs (Iyengar 2005; Kaufman and Schaefer 2013), researchers have sought to generalize the ideas of modified policy iteration from MDPs to MGs and RMDPs. However, the attempts to speed up policy iteration while guaranteeing convergence have been largely unsuccessful (Pérolat et al. 2016). Existing algorithms are either too slow for larger problems or lack optimality guarantees.
Policy iteration (Puterman 2005), another basic MDP algorithm, can dramatically reduce the number of Bellman operator evaluations and compute the optimal policy in strongly polynomial time in MDPs (Ye 2011). Hoffman-Karp algorithm, also known as robust policy iteration (Iyengar 2005), for MGs and RMDPs adapts policy iteration to MGs and RMDPs. Although Hoffman-Karp has polynomial worst-case time complexity (Hansen, Miltersen, and Zwick 2013), it can be slower than value iteration in practice. Each Hoffman-Karp policy evaluation requires computing the adversarial agent’s optimal policy. That is a significant increase in the complexity of the policy evaluation step in MDPs, which entails solving a system of linear equations.
@@
One well-known attempt to fix PAI’s non-convergence is the Filar-Tolwinski (FT) algorithm (Filar and Tolwinski 1991). It leverages the observation that PAI can be seen as Newton’s method on the $L_{2}$ norm of the Bellman residual. FT replaces the pure Newton’s method of PAI with the modified Newton’s method, which uses Armijo’s rule when deciding the step size in the value function update. While (Filar and Tolwinski 1991) claims that this resolves the cycling issues found with PAI, we show that FT may not converge. We discuss this issue in more detail in Section˜4.
Recent years have seen several notable attempts to develop algorithms that match the empirical performance of PAI while guaranteeing convergence to an optimal policy. Robust Modified Policy Iteration (RMPI) (Kaufman and Schaefer 2013) and Partial Policy Iteration (PPI) (Ho, Petrik, and Wiesemann 2021) modify Hoffman-Karp to evaluate the adversarial policy approximately. RMPI uses a fixed-precision approximation, while PPI adapts the evaluation throughout the algorithm’s execution. Numerical evidence suggests that PPI outperforms RMPI (Ho, Petrik, and Wiesemann 2021). The Winnicki-Srikant (WS) algorithm combines value iteration steps with policy backup steps and proposes ratios that guarantee the algorithm’s convergence (Winnicki and Srikant 2023).
@@ 3.2 Markov Games
Markov games extend Markov decision processes to a zero-sum game-theoretic setting (Kallenberg 2022; Filar and Vrieze 1997) and can model multi-agent reinforcement learning. An imperfect-information Markov game is defined as $(\mathcal{S},\mathcal{A},\mathcal{B},r,P,s_{0})$ where $\mathcal{S}=\left\{1,\dots,S\right\}$ is the finite non-empty set of states that the agents share, $\mathcal{A}=\left\{1,\dots,A\right\}$ is the finite non-empty set of actions for the primary agent, $\mathcal{B}=\left\{1,\dots,B\right\}$ is the finite non-empty set of actions for the adversarial agent. The function $r\colon\mathcal{S}\times\mathcal{A}\times\mathcal{B}\times\mathcal{S}\to[-r_{\max},r_{\max}]$ for $r_{\max}\in\mathbb{R}$ represents the rewards the primary agent seeks to maximize and the adversarial agent seeks to minimize. The function $P\colon\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\Delta^{\mathcal{S}}$ is the transition probability function, where $p(s,a,b,s^{\prime})$ is the probability of transitioning from state $s$ to state $s^{\prime}$ after the agents take actions action $a$ and $b$, respectively. Finally, $s_{0}\in\mathcal{S}$ is the initial state.
We consider infinite-horizon discounted rewards for a discount factor $\gamma\in(0,1)$, and restrict attention to randomized stationary policies $\Pi:=(\Delta^{\mathcal{A}})^{\mathcal{S}}$ and $\Sigma:=(\Delta^{\mathcal{B}})^{\mathcal{S}}$ for the maximizing and minimizing agents, respectively. Note that the restriction to randomized stationary policies is not limiting, because neither of the players can benefit from using Markov or history-dependent policies (Kallenberg 2022; Filar and Vrieze 1997). The value function $\bm{v}^{\bm{\pi},\bm{\sigma}}\in\mathbb{R}^{S}$ associated with each $\bm{\pi}\in\Pi$ and $\bm{\sigma}\in\Sigma$ as (Filar and Vrieze 1997):
$v^{\bm{\pi},\bm{\sigma}}_{s}\;:=\;\mathbb{E}_{\bm{\pi},\bm{\sigma}}^{s}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\tilde{s}_{t},\tilde{a}_{t},\tilde{b}_{t},\tilde{s}_{t+1})\right],\quad\forall s\in\mathcal{S}.$ (2)
@@
$(\bm{\pi^{\star}},\bm{\sigma^{\star}})\in\mathfrak{S}_{\epsilon}(\rho),\quad\text{where}\quad\rho_{\mathrm{G}}(s_{0},\bm{\pi},\bm{\sigma})\;:=\;v_{s_{0}}^{\bm{\pi},\bm{\sigma}}.$ (4)
@@
$\mathfrak{T}^{\bm{\pi},\bm{\sigma}}_{s}\bm{v}\;:=\;r_{s}^{\bm{\pi},\bm{\sigma}}+\gamma\cdot\smashoperator[]{\sum_{(a,b)\in\mathcal{A}_{s}\times\mathcal{B}_{s}}^{}}\pi_{a}(s)\cdot\sigma_{b}(s)\cdot(\bm{p}^{\bm{\pi},\bm{\sigma}}_{s})^{\mathsf{T}}\bm{v}~.$ (5)
@@
The set $\mathfrak{B}^{\star}\bm{v}$ is non-empty for each $\bm{v}\in\mathbb{R}^{S}$, and
$\mathfrak{B}^{\star}\bm{v}\;\subseteq\;\mathfrak{S}_{\epsilon}(\rho_{\mathrm{G}})\quad\text{where}\quad\epsilon=\frac{2\gamma}{1-\gamma}\psi_{\infty}(\bm{v}).$
The proof, which we include in the appendix for the sake of completeness, follows standard arguments; see, for example, (Kallenberg 2022, Theorem 10.11). We note that the bound in Proposition˜3.1 is tighter than the bounds given, for example, in Ho, Petrik, and Wiesemann (2021, Corollary A.4) and Williams and Baird (1993, Theorems 3.1, 3.2).
@@ 3.3 Robust MDPs
RMDPs generalize MDPs to allow for adversarial perturbations to the transition probabilities. We consider s-rectangular RMDPs $(\mathcal{S},\mathcal{A},r,\mathcal{P},s_{0})$ where $\mathcal{S}$ and $\mathcal{A}$ are the finite non-empty sets of states and actions, respectively (Wiesemann, Kuhn, and Rustem 2013; Ho, Petrik, and Wiesemann 2021), and $r\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[-r_{\max},r_{\max}]$ is the reward function. The ambiguity set $\mathcal{P}:=(\mathcal{P}_{s})_{s\in\mathcal{S}}$, where $\mathcal{P}_{s}\subseteq\Delta^{\mathcal{S}}$ is compact and non-empty for each $s\in\mathcal{S}$, and determines the range of possible adversarial transition probability functions. Finally, the initial state is $s_{0}$.
It is common to define the ambiguity sets in RMDPs as bounded norm-balls around a given nominal transition function $\bm{\bar{p}}\colon\mathcal{S}\times\mathcal{A}\to\Delta^{S}$, such as (Ho, Petrik, and Wiesemann 2021, 2022; Behzadian et al. 2021; Behzadian, Petrik, and Ho 2021)
@@
$\begin{gathered}\max_{\bm{\pi}\in\Pi}\min_{\bm{p}\in\mathcal{P}}\,\rho_{\mathrm{R}}(\bm{\pi},\bm{p}),\\ \rho_{\mathrm{R}}(\bm{\pi},\bm{p}):=\mathbb{E}_{\bm{\pi},\bm{p}}^{s_{0}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\tilde{s}_{t},\tilde{a}_{t},\tilde{s}_{t+1})\right],\end{gathered}$ (8)
where $\gamma\in(0,1)$. We emphasize that for each $\bm{\pi}\in\Pi$, the domain of $\rho_{\mathrm{R}}(\bm{\pi},\cdot)$ is the set of feasible transition probabilities $\mathcal{P}$, rather than the set of all transition probabilities. An optimal policy $\bm{\pi^{\star}}$ in (8) exists and can be computed by robust value or policy iteration (Wiesemann, Kuhn, and Rustem 2013; Iyengar 2005).
@@
For RMDPs, value functions and Bellman operators are defined analogously to how they are defined for MGs. We summarize these definitions in Section˜A.2. In the remainder of the paper, we describe the algorithms for MGs which generalize readily to RMDPs.
Computationally, the main difference between RMDPs and MGs is in computing the Bellman operator. For most common ambiguity sets $\mathcal{P}$, the robust Bellman operator can be implemented by solving a convex optimization problem (Wiesemann, Kuhn, and Rustem 2013). For the $L_{1}$-bound ambiguity sets, the robust Bellman operator can be implemented by solving a linear program (Ho, Petrik, and Wiesemann 2021, Appendix C). Significantly more efficient methods exist for ambiguity sets bounded by norms and $\varphi$-divergences (Ho, Petrik, and Wiesemann 2021, 2022; Behzadian, Petrik, and Ho 2021).
4 Filar-Tolwinski Algorithm Does Not Converge
Pollatschek and Avi-Itzhak (Pollatschek and Avi-Itzhak 1969) proposed one of the first alternatives to value iteration (Shapley 1953) for solving MGs. This algorithm, which we refer to as the PAI algorithm, can be viewed as applying Newton’s method to the problem of finding a zero of $\psi_{2}(\bm{v})^{2}$. While PAI is known to converge to the optimal value function $\bm{v}^{\star}$ under certain restrictive conditions (Pollatschek and Avi-Itzhak 1969, Theorem 5), it is also known to not converge at all for certain MGs (Van der Wal 1978). Filar and Tolwinski (Filar and Tolwinski 1991) proposed a modified Newton method intended to fix this convergence issue. In this section, we provide a counterexample to (Filar and Tolwinski 1991, Theorem 3.3), where it is claimed that the modified Newton method converges from some constant initial vector to $\bm{v}^{\star}$. The Filar-Tolwinski (FT) algorithm is described in Algorithm˜1.

### v2

@@ 1 Introduction
Markov Games (MG) (Kallenberg2022) and Robust MDPs (RMDPs) (Iyengar2005; Wiesemann2013; Ho2022) are two important models that generalize Markov Decision Processes (MDPs) (Puterman2005). Markov games can model strategic adversaries that can act to minimize the agent’s returns and are a common model in multi-agent reinforcement learning (Shou2022; Littman1994a). Similarly, RMDPs can model an adversarial nature that can perturb transition probabilities and rewards to minimize the agent’s returns and are useful when making decisions with imperfect data-driven models (Lobo2023; Behzadian2021). In recent years, MGs and RMDPs have seen an increasing number of applications in machine and reinforcement learning, which has motivated the study of efficient algorithms for solving them (Perolat2016; Ho2021; Ho2022; Behzadian2021a; Kaufman2013; Winnicki2023).
@@
Although basic algorithms, like value and policy iteration, adapt readily from MDPs to MGs and RMDPs, developing more efficient algorithms has been challenging. The efforts to adapt efficient optimistic policy iteration (OPI) algorithms, such as modified or fitted policy iteration, have been difficult. Many natural OPI algorithms proposed for MGs and RMDPs cycle among suboptimal policies, alternatively improving the minimization or maximization sides of the saddle point equilibrium. The lack of convergence is often counter-intuitive and has led to several incorrect convergence proofs in the literature (Condon1993; Perolat2016; Filar1991). The overarching reason is that OPI in MG and RMDPs do not monotonically improve the policy and its value function as in MDPs.
@@
We make two main contributions in this paper. First, we show that a fast OPI method proposed in Filar1991 can terminate with an arbitrarily suboptimal policy. Perolat2016 first identified a gap in the proof of correctness in Filar1991 but hypothesized the algorithm works nevertheless. In contrast, we show that the algorithm is inherently suboptimal.
@@
Second, we propose and analyze Residual Conditioned Policy Iteration (RCPI). RCPI is a new, simple approximate policy iteration algorithm for solving MGs and RMDPs that is guaranteed to converge to optimal policies. It builds on earlier efficient OPI algorithms (Filar1991; Perolat2016; Ho2021; Winnicki2023) and combines them with an adaptive correction step. Our theoretical analysis shows that RCPI matches the worst-case computational complexity of value iteration. Our numerical results show that on a wide range of problems, RCPI outperforms other convergent algorithms by several orders of magnitude, even in moderately sized problems.
@@
In this paper, we restrict our focus to model-based algorithms for MGs and RMDPs. It is important to note that this setting differs from online algorithms for solving games and multi-agent reinforcement learning problems, such as in Zhang2022. Although some of the issues that need to be overcome in online and model-based solvers are similar, we leave the study of the exact relationship between online and model-based algorithms for future work.
@@
The remainder of the paper is organized as follows. Section˜2 positions our work in the context of prior algorithmic developments for MGs and RMDPs. Then, Section˜3 describes the formal framework for MGs and RMDPs. Section˜4 describes our first contribution, which is to show that an existing OPI algorithm (Filar1991) may fail with an arbitrarily suboptimal policy. Section˜5 describes our second and main contribution, the RCPI algorithm, along with its convergence rate and computational complexity analysis. Finally, our numerical results in Section˜6 compare RCPI with existings algorithms for solving MGs and RMDPs.
@@ 2 Prior Work: Solving MGs and RMDPs
In this section, we summarize prior efforts on developing OPI algorithms for MGs and RMDPs. We note that MG and RMDP communities have been largely separate, though the similarities between them have been noted and exploited previously (Iyengar2005; Grand-Clement2024; Grand-Clement2025).
Value iteration is a simple convergent algorithm for solving MDPs, RMDPs, and MGs, but can be very slow in many practical settings (Puterman2005). Many faster convergent algorithms for MDPs exist, such as policy iteration or modified policy iteration. Since the early days of MG (Condon1993) and RMDPs (Iyengar2005; Kaufman2013), researchers have sought to generalize the ideas of modified policy iteration from MDPs to MGs and RMDPs. However, the attempts to speed up policy iteration while guaranteeing convergence have been largely unsuccessful (Perolat2016). Existing algorithms are either too slow for larger problems or lack optimality guarantees.
Policy iteration (Puterman2005), another basic MDP algorithm, can dramatically reduce the number of Bellman operator evaluations and compute the optimal policy in strongly polynomial time in MDPs (Ye2011). Hoffman-Karp algorithm, also known as robust policy iteration (Iyengar2005), for MGs and RMDPs adapts policy iteration to MGs and RMDPs. Although Hoffman-Karp has polynomial worst-case time complexity (Hansen2013), it can be slower than value iteration in practice. Each Hoffman-Karp policy evaluation requires computing the adversarial agent’s optimal policy. That is a significant increase in the complexity of the policy evaluation step in MDPs, which entails solving a system of linear equations.
@@
One well-known attempt to fix PAI’s non-convergence is the Filar-Tolwinski (FT) algorithm (Filar1991). It leverages the observation that PAI can be seen as Newton’s method on the $L_{2}$ norm of the Bellman residual. FT replaces the pure Newton’s method of PAI with the modified Newton’s method, which uses Armijo’s rule when deciding the step size in the value function update. While (Filar1991) claims that this resolves the cycling issues found with PAI, we show that FT may not converge. We discuss this issue in more detail in Section˜4.
Recent years have seen several notable attempts to develop algorithms that match the empirical performance of PAI while guaranteeing convergence to an optimal policy. Robust Modified Policy Iteration (RMPI) (Kaufman2013) and Partial Policy Iteration (PPI) (Ho2021) modify Hoffman-Karp to evaluate the adversarial policy approximately. RMPI uses a fixed-precision approximation, while PPI adapts the evaluation throughout the algorithm’s execution. Numerical evidence suggests that PPI outperforms RMPI (Ho2021). The Winnicki-Srikant (WS) algorithm combines value iteration steps with policy backup steps and proposes ratios that guarantee the algorithm’s convergence (Winnicki2023).
@@ 3.2 Markov Games
Markov games extend Markov decision processes to a zero-sum game-theoretic setting (Kallenberg2022; Filar1996) and can model multi-agent reinforcement learning. An imperfect-information Markov game is defined as $(\mathcal{S},\mathcal{A},\mathcal{B},r,P,s_{0})$ where $\mathcal{S}=\left\{1,\dots,S\right\}$ is the finite non-empty set of states that the agents share, $\mathcal{A}=\left\{1,\dots,A\right\}$ is the finite non-empty set of actions for the primary agent, $\mathcal{B}=\left\{1,\dots,B\right\}$ is the finite non-empty set of actions for the adversarial agent. The function $r\colon\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to[-r_{\max},r_{\max}]$ for $r_{\max}\in\mathbb{R}$ represents the rewards the primary agent seeks to maximize and the adversarial agent seeks to minimize. The function $P\colon\mathcal{S}\times\mathcal{A}\times\mathcal{B}\to\Delta^{\mathcal{S}}$ is the transition probability function, where $p(s,a,b,s^{\prime})$ is the probability of transitioning from state $s$ to state $s^{\prime}$ after the agents take actions action $a$ and $b$, respectively. Finally, $s_{0}\in\mathcal{S}$ is the initial state.
We consider infinite-horizon discounted rewards for a discount factor $\gamma\in(0,1)$, and restrict attention to randomized stationary policies $\Pi:=(\Delta^{\mathcal{A}})^{\mathcal{S}}$ and $\Sigma:=(\Delta^{\mathcal{B}})^{\mathcal{S}}$ for the maximizing and minimizing agents, respectively. Note that the restriction to randomized stationary policies is not limiting, because neither of the players can benefit from using Markov or history-dependent policies (Kallenberg2022; Filar1996). The value function $\bm{v}^{\bm{\pi},\bm{\sigma}}\in\mathbb{R}^{S}$ associated with each $\bm{\pi}\in\Pi$ and $\bm{\sigma}\in\Sigma$ as (Filar1996):
$v^{\bm{\pi},\bm{\sigma}}_{s}\;:=\;\mathbb{E}_{\bm{\pi},\bm{\sigma}}^{s}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\tilde{s}_{t},\tilde{a}_{t},\tilde{b}_{t})\right],\quad\forall s\in\mathcal{S}.$ (2)
@@
$(\bm{\pi^{\star}},\bm{\sigma^{\star}})\in\mathfrak{S}_{\epsilon}(\rho),\;\text{where}\;\rho_{\mathrm{G}}(s_{0},\bm{\pi},\bm{\sigma})\;:=\;v_{s_{0}}^{\bm{\pi},\bm{\sigma}}.$ (4)
@@
$\mathfrak{T}^{\bm{\pi},\bm{\sigma}}_{s}\bm{v}\;:=\;r_{s}^{\bm{\pi},\bm{\sigma}}+\gamma\cdot\bm{P}^{\bm{\pi},\bm{\sigma}}_{s}\bm{v}~.$ (5)
@@
For each $\bm{v}\in\mathbb{R}^{S}$:
$\emptyset\neq\mathfrak{B}^{\star}\bm{v}\;\subseteq\;\mathfrak{S}_{\epsilon}(\rho_{\mathrm{G}}),\quad\text{where}\quad\epsilon=\frac{2\gamma}{1-\gamma}\psi_{\infty}(\bm{v}).$
The proof, which we include in the appendix for the sake of completeness, follows standard arguments; see, for example, (Kallenberg2022, theorem 10.11). We note that the bound in Proposition˜3.1 is tighter than the bounds given, for example, in Ho2021 and Williams1993a.
@@ 3.3 Robust MDPs
RMDPs generalize MDPs to allow for adversarial perturbations to the transition probabilities. We consider s-rectangular RMDPs $(\mathcal{S},\mathcal{A},r,\mathcal{P},s_{0})$ where $\mathcal{S}$ and $\mathcal{A}$ are the finite non-empty sets of states and actions, respectively (Wiesemann2013; Ho2021), and $r\colon\mathcal{S}\times\mathcal{A}\to[-r_{\max},r_{\max}]$ is the reward function. The ambiguity set $\mathcal{P}:=(\mathcal{P}_{s})_{s\in\mathcal{S}}$, where $\mathcal{P}_{s}\subseteq\Delta^{\mathcal{S}}$ is compact and non-empty for each $s\in\mathcal{S}$, and determines the range of possible adversarial transition probability functions. Finally, the initial state is $s_{0}$.
It is common to define the ambiguity sets in RMDPs as bounded norm-balls around a given nominal transition function $\bm{\bar{p}}\colon\mathcal{S}\times\mathcal{A}\to\Delta^{S}$, such as (Ho2021; Ho2022; Behzadian2021; Behzadian2021a)
@@
$\begin{gathered}\max_{\bm{\pi}\in\Pi}\min_{\bm{p}\in\mathcal{P}}\,\rho_{\mathrm{R}}(\bm{\pi},\bm{p}),\\ \rho_{\mathrm{R}}(\bm{\pi},\bm{p}):=\mathbb{E}_{\bm{\pi},\bm{p}}^{s_{0}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\tilde{s}_{t},\tilde{a}_{t})\right],\end{gathered}$ (8)
where $\gamma\in(0,1)$. We emphasize that for each $\bm{\pi}\in\Pi$, the domain of $\rho_{\mathrm{R}}(\bm{\pi},\cdot)$ is the set of feasible transition probabilities $\mathcal{P}$, rather than the set of all transition probabilities. An optimal policy $\bm{\pi^{\star}}$ in (8) exists and can be computed by robust value or policy iteration (Wiesemann2013; Iyengar2005).
@@
For RMDPs, value functions and Bellman operators are defined analogously to how they are defined for MGs. For more detail, please see the appendix. In the remainder of the paper, we describe the algorithms for MGs which generalize to RMDPs.
Computationally, the main difference between RMDPs and MGs is in computing the Bellman operator. For most common ambiguity sets $\mathcal{P}$, the robust Bellman operator can be implemented by solving a convex optimization problem (Wiesemann2013). For the $L_{1}$-bound ambiguity sets, the robust Bellman operator can be implemented by solving a linear program (Ho2021, appendix C). Significantly more efficient methods exist for ambiguity sets bounded by norms and $\varphi$-divergences (Ho2021; Ho2022; Behzadian2021a).
4 Filar-Tolwinski Algorithm
May Not Converge
PAI1969 proposed one of the first alternatives to value iteration (Shapley1953) for solving MGs. This algorithm, which we refer to as the PAI algorithm, can be viewed as applying Newton’s method to the problem of finding a zero of $\psi_{2}(\bm{v})^{2}$. While PAI is known to converge to the optimal value function $\bm{v}^{\star}$ under certain restrictive conditions (PAI1969, theorem 5), it is also known to not converge at all for certain MGs (VanderWal1978). Filar and Tolwinski (Filar1991) proposed a modified Newton method intended to fix this convergence issue. In this section, we provide a counterexample to Filar1991, where it is claimed that the modified Newton method converges from some constant initial vector to $\bm{v}^{\star}$. The Filar-Tolwinski (FT) algorithm is described in Algorithm˜1.